# Trustworthy Transfer Learning: A Survey

### v1

@@ 3.1.1 Distribution Discrepancy
    where $\phi(\cdot)$ is a convex lower semi-continuous function that satisfies $\phi(1)=0$. $\phi^{*}$ is the conjugate function of $\phi$. $\mathcal{T}$ is a set of measurable functions. Based on this observation, ? (?) define a notion of $f$-divergence guided discrepancy as follows.
    $d_{\phi}(\mathcal{D}_{S},\mathcal{D}_{T})=\sup_{h^{\prime}\in\mathcal{H}}\left|\mathbb{E}_{x\sim\mathrm{P}_{S}}\left[\ell\left(h(x),h^{\prime}(x)\right)\right]-\mathbb{E}_{x\sim\mathrm{P}_{T}}\left[\phi^{*}\left(\ell\left(h(x),h^{\prime}(x)\right)\right)\right]\right|$ (7)
  • Generalized Discrepancy (?): In contrast, ? (?) generalize the discrepancy distance in Eq. (4) using reweighting techniques. That is, the difference between two distributions can be adjusted by multiplying the loss for each training example by a non-negative weight (?, ?, ?). Formally, for any hypothesis-dependent reweighting function $U_{h}$, the generalized discrepancy is defined as follows.
    $d_{\mathrm{DISC}}(\mathcal{D}_{S},\mathcal{D}_{T})=\max_{h\in\mathcal{H},h^{\prime\prime}\in\mathcal{H}^{\prime\prime}}\left|\mathbb{E}_{x\sim\hat{\mathrm{P}}_{S}}\left[{U}_{h}(x)\cdot\ell\left(h\left(x\right),f_{S}\left(x\right)\right)\right]-\mathbb{E}_{x\sim\hat{\mathrm{P}}_{T}}\left[\ell\left(h(x),h^{\prime\prime}(x)\right)\right]\right|$ (8)
    where $\hat{\mathrm{P}}_{S},\hat{\mathrm{P}}_{T}$ denotes the empirical distributions of source and target domains, respectively, and $f_{S}$ denotes the source labeling function.
  • Rényi Divergence (?, ?): Furthermore, ? (?) and ? (?) derive the generalization bounds for adaptation approaches based on importance reweighting (e.g., sample reweighting for single-source adaptation (?) and domain reweighting for multi-source adaptation (?)) based on the following Rényi divergence (?).
    $d_{\alpha}(\mathcal{D}_{T}||\mathcal{D}_{S})=\frac{1}{\alpha-1}\log\sum_{x\in\mathcal{X}}\mathrm{P}_{T}(x)\left(\frac{\mathrm{P}_{T}(x)}{\mathrm{P}_{S}(x)}\right)^{\alpha-1}$ (9)
    where $\alpha\geq 0$.
  • Wasserstein Distance (?): In general, fro any $p\geq 1$, the $p$-Wasserstein distance between two distributions can be defined as follows.
@@
    where $\Pi(\mathrm{P}_{S},\mathrm{P}_{T})$ is the set of all measures over $\mathcal{X}\times\mathcal{X}$ with marginals $\mathrm{P}_{S}$ and $\mathrm{P}_{T}$, and $c(\cdot,\cdot)$ is a distance function. When $p=1$, 1-Wasserstein distance (also known as earth mover’s distance) is one special case of the integral probability metric with =$\mathcal{H}=\{h:||h||_{L}\leq 1\}$. Specifically, based on the Kantorovich-Rubinstein duality (?), it holds that
[…]
    where $||h||_{L}=\sup_{x\neq x^{\prime}}|h(x)-h(x^{\prime})|/c(x,x^{\prime})$. This enables a practical empirical estimation of the 1-Wasserstein distance using gradient descent optimization (?). Therefore, ? (?) and ? (?) apply the 1-Wasserstein distance to analyze distribution shifts between source and target domains.
  • Maximum Mean Discrepancy (?, ?): ? (?) and ? (?) leverage the maximum discrepancy discrepancy (MMD) (?) to measure the distribution difference between source and target domains. MMD can be considered as another special case of the integral probability metric by instantiating the hypothesis space with a unit ball in a reproducing kernel Hilbert space associated with kernel $k(\cdot,\cdot)$. Given a kernel function $k(\cdot,\cdot)$, the MMD between source and target distribution can be defined as:
    $d_{\mathrm{MMD}}(\mathcal{D}_{S},\mathcal{D}_{T})$ $=\mathbb{E}_{x_{S},x^{\prime}_{S}\sim\mathrm{P}_{S}}\big{[}k\left(x_{S},x^{\prime}_{S}\right)\big{]}-2\mathbb{E}_{x_{S}\sim\mathrm{P}_{S},x_{T}\sim\mathrm{P}_{T}}\big{[}k\left(x_{S},x_{T}\right)\big{]}$ (12)
@@
    ? (?) further show the generalization error bounds based on MMD.
@@
  • Cauchy-Schwarz Divergence (?): Recently, ? (?) use the Cauchy-Schwarz (CS) divergence (?) to theoretically understand the knowledge transferability across domains.
[…]
@@
    It is shown (?) that this CS divergence can lead to tighter generalization error bounds than KL divergence (?). Besides, the empirical estimate of the CS divergence is closely related to MMD (?).
@@
From the perspective of empirical estimation, the discrepancy measures can be broadly categorized into two groups. The first group includes statistical discrepancy measures (?, ?, ?), such as Maximum Mean Discrepancy (MMD) (?, ?) and Wasserstein distance (?, ?, ?), which can be directly estimated from finite samples. The second group is based on adversarial learning (?, ?, ?, ?, ?, ?), which requires an additional neural network to optimize an adversarial objective. More recently, ? (?) and ? (?) provide empirical comparisons of various discrepancy measures in natural language processing and computer vision tasks. It is noted that when there are no labeled samples in the target domain, one common strategy in designing practical domain adaptation algorithms is to minimize the discrepancy across domains over $\mathcal{X}$. However, it has been shown (?, ?, ?, ?) that exact marginal distribution matching might lead to negative transfer in practice.
@@
The notion of distribution discrepancy has been applied to understand knowledge transferability in various realistic adaptation scenarios, including single-source adaptation (?, ?, ?, ?, ?), multi-source adaptation (?, ?, ?), open-set adaptation (?, ?), domain generalization (?, ?) (also known as out-of-distribution generalization), privacy-preserving federated adaptation (?), dynamic adaptation (?, ?), etc.
3.1.2 Task Diversity
@@
Task diversity (?, ?) is another tool for theoretically understanding the performance of transfer learning. It enables a relaxed data assumption that the source and target domains can have different output spaces, i.e., each domain can be associated with a different learning task (?). In the context of transfer learning, it assumes that a generic nonlinear feature representation function is shared across all tasks. Then each task is associated with a shared representation learning function and a task-specific prediction function. In (?), task diversity is defined to characterize how the worst-case representation difference can be controlled when the task-averaged representation difference is small. In this case, the worst-case representation difference is the distance between two representation functions with the worst-case task-specific prediction function, while the task-averaged representation difference indicates the distance between two representation functions over all the training tasks.
Definition 3 (Task Diversity (?))
@@
$\underbrace{\sup_{h_{0}\in\mathcal{H}_{0}}\inf_{h^{\prime}\in\mathcal{H}}\left\{\mathcal{E}_{T}\left(h^{\prime}\circ g^{\prime}\right)-\mathcal{E}_{T}\left(h_{0}\circ g\right)\right\}}_{\text{Worst-case representation difference}}\leq\frac{1}{\nu}\cdot\underbrace{\left(\frac{1}{N}\sum_{i=1}^{N}\inf_{h^{\prime}\in\mathcal{H}}\left\{\mathcal{E}_{S_{i}}\left(h^{\prime}\circ g^{\prime}\right)-\mathcal{E}_{S_{i}}\left(h_{i}\circ g\right)\right\}\right)}_{\text{Task-averaged representation difference}}~{}~{}+~{}~{}~{}\epsilon$ (14)
@@
Based on the task diversity, ? (?) derive the excess risk bounds of transfer learning for the target task in terms of the complexity of the shared representation function class $\mathcal{G}$, the complexity of the prediction function class $\mathcal{H}$, the number of tasks $N$, and the number of training samples for each task (both source and target). Furthermore, ? (?) show that under the Lipschitz assumption for the loss function, the excess risk in the target task only achieves the standard rate of $\mathcal{O}(n_{T}^{-1/2})$, where $n_{T}$ is the number of training samples in the target task. By using the smoothness assumption for the loss function (?), they derive optimistic rates that interpolate between the standard rate of $\mathcal{O}(n_{T}^{-1/2})$ and the fast rate of $\mathcal{O}(n_{T}^{-1})$ for the excess risk in the target task.
@@
In addition, ? (?) and ? (?) consider a simplified version of task diversity in the cases of linear prediction functions and quadratic loss. They also theoretically show the benefits (i.e., reduced sample complexity in the target task induced by all available source samples) of representation learning from source tasks. Based on the task diversity, ? (?) analyze more realistic learning scenarios in which the source and target tasks use different prediction function spaces. However, all the aforementioned theoretical analyses assume uniform sampling from each source task, i.e., all source tasks are equally important for learning a representation function. Instead, ? (?, ?) and ? (?) study active transfer learning by quantifying the task relatedness and selecting the source tasks that are most relevant to the target task. Similarly, ? (?) explore the selection of source tasks for multi-task fine-tuning of foundation models, e.g., fine-tuning the foundation model on auxiliary source tasks before adapting it to the target task with limited labeled samples. More recently, ? (?) show that pre-training on a single source task with a high diversity of classes can provably improve the sample efficiency of the downstream tasks. In contrast, ? (?) leverage task diversity to understand the in-context learning behavior of foundation models.
3.1.3 Transferability Estimation
@@
In contrast to the data-centric transferability analyses in Subsection 3.1.1 and Subsection 3.1.2, this subsection explores the knowledge transferability of pre-trained source models. This is driven by the rapidly expanding open-source model repositories such as Hugging Face (?) and PyTorch Hub (?). Fine-tuning a pre-trained source model on downstream target data sets with limited sample sizes improves model accuracy and robustness (?). A natural question arises in this scenario: Given a large pool of pre-trained source models, how can we efficiently select the best one for a target data set? As shown in Figure 3, another relevant question is how to identify the most suitable domains/tasks for a given pre-trained source model. One trivial solution is brute-force fine-tuning, where all source models are fine-tuned individually and then ranked based on their transfer accuracy. However, this method is highly time-consuming and computationally expensive. To solve this problem, transferability estimation has been studied to quantitatively measure how effectively the knowledge can be transferred from a pre-trained source model to a target domain (?, ?, ?, ?, ?). Following (?), given a pre-trained source model $f_{S}(\cdot)$, the transferability from $f_{S}(\cdot)$ to a target domain associated with sampling distribution $\mathrm{P}_{T}$ can be defined below.
Definition 4 (Transferability Measure (?))
@@
$\mathrm{Trf}(S\to T)=\mathbb{E}_{(x,y)\sim\mathrm{P}_{T}}\left[\mathrm{acc}\left(x,y;f_{T}\right)\right]$ (15)
@@
Figure 3: Evaluation of transferability between the pre-trained source model and the target data: (a) Transferability scores select the best source model for the given target data given a large pool of pre-trained source models. (b) Transferability scores identify the most suitable application domains/tasks for a source model.
@@
The estimation of the transferability score is related to the architectures of pre-trained source models. In scenarios where the source models are trained on supervised classification tasks, every source model $f_{S}(\cdot)$ is associated with a feature extractor and a predictor. In this case, NCE (?) leverages conditional entropy to define the transferability score, assuming that source and target domains share the same input samples but different labels. It is motivated by the observation that the optimal average log-likelihood on target training samples is lower bounded by the negative conditional entropy. Similarly, LEEP (?) estimates the average log-likelihood of target samples using the dummy label distributions generated from the pre-trained source model. Notably, it is shown that LEEP is an upper bound of the NCE (?) plus the average log-likelihood of the dummy labels. It is computationally efficient by using only a single forward pass of the source model through the target data. Nevertheless, LEEP can not handle unsupervised and self-supervised pre-trained models with only a feature extractor.
[…]
@@
The transferability metrics mentioned above enable the selection of the best source model from a large pool of open-sourced pre-trained models. Recent studies (?, ?) take one step further by studying source model ensemble selections, which transfer knowledge from multiple pre-trained source models to target training samples. This line of research is inspired by the success of ensemble machine learning models (?) in improving model performance. Specifically, ? (?) extends LEEP (?) to select source model ensembles under the assumption that the source models operate independently. In contrast, OSBORN (?) relaxes this assumption and explores the inter-model cohesion among source models to estimate the transferability of an ensemble of models to a target domain.
3.2 Non-IID Transferability

### v2

@@ 3.1.1. Distribution Discrepancy
    where $h||_{L}=\sup_{x\neq x^{\prime}}|h(x)-h(x^{\prime})|/c(x,x^{\prime})$. This enables a practical empirical estimation of the 1-Wasserstein distance using gradient descent optimization (arjovsky2017wasserstein). Therefore, (Wasserstein; redko2017theoretical) apply the 1-Wasserstein distance to analyze distribution shifts between source and target domains.
  • Maximum Mean Discrepancy: (tzeng2014deep; long2015learning) leverage the maximum discrepancy discrepancy (MMD) (MMD) to measure the distribution difference between source and target domains. MMD can be considered as another special case of the integral probability metric by instantiating the hypothesis space with a unit ball in a reproducing kernel Hilbert space associated with kernel $k(\cdot,\cdot)$. Given a kernel function $k(\cdot,\cdot)$, the MMD between source and target distribution can be defined as:
    (12) $d_{\mathrm{MMD}}(\mathcal{D}_{S},\mathcal{D}_{T})$ $=\mathbb{E}_{x_{S},x^{\prime}_{S}\sim\mathrm{P}_{S}}\big[k\left(x_{S},x^{\prime}_{S}\right)\big]-2\mathbb{E}_{x_{S}\sim\mathrm{P}_{S},x_{T}\sim\mathrm{P}_{T}}\big[k\left(x_{S},x_{T}\right)\big]$
@@
    (redko2019advances) further show the generalization error bounds based on MMD.
@@
  • Cauchy-Schwarz Divergence (yin2024domain): Recently, (yin2024domain) use the Cauchy-Schwarz (CS) divergence (principe2010information) to theoretically understand the knowledge transferability across domains.
    (13) $d_{\mathrm{CS}}(\mathcal{D}_{S},\mathcal{D}_{T})=-\log\left(\frac{\left(\int p_{S}(x)p_{T}(x)dx\right)^{2}}{\int p_{S}^{2}(x)dx\cdot\int p_{T}^{2}(x)dx}\right)$
@@
    It is shown (yin2024domain) that this CS divergence can lead to tighter generalization error bounds than the KL divergence (nguyen2022kl). Besides, the empirical estimate of the CS divergence is closely related to MMD (MMD).
@@
From the perspective of empirical estimation, the discrepancy measures can be broadly categorized into two groups. The first group includes statistical discrepancy measures (nguyen2022kl; CORAL; RSD), such as Maximum Mean Discrepancy (MMD) (tzeng2014deep; long2015learning) and Wasserstein distance (courty2016optimal; redko2017theoretical; fatras2021unbalanced), which can be directly estimated from finite samples. The second group is based on adversarial learning (DANN; saito2018maximum; tzeng2017adversarial; hoffman2018cycada; MDD; f_divergence), which requires an additional neural network to optimize an adversarial objective. More recently, (KashyapHKZ21; YuanHWMLSH22) provide empirical comparisons of various discrepancy measures in natural language processing and computer vision tasks. It is noted that when there are no labeled samples in the target domain, one common strategy in designing practical domain adaptation algorithms is to minimize the discrepancy across domains over $\mathcal{X}$. However, it has been shown (Ben-DavidLLP10; 0002CZG19; WuWKL19; JohanssonSR19) that exact marginal distribution matching might lead to negative transfer in practice.
@@
The notion of distribution discrepancy has been applied to understand knowledge transferability in various realistic adaptation scenarios, including single-source adaptation (cortes2010learning; A_distance; MDD; f_divergence; nguyen2022kl), multi-source adaptation (mansour2021theory; hoffman2018algorithms; wu2024distributional), open-set adaptation (fang2020open; RAINCOAT), domain generalization (MuandetBS13; BlanchardLS11) (also known as out-of-distribution generalization), privacy-preserving federated adaptation (PengHZS20), dynamic adaptation (KumarML20; WuH22), etc.
3.1.2. Task Diversity
@@
Task diversity (task_diversity; WatkinsUNA23) is another tool for theoretically understanding the performance of transfer learning. It enables a relaxed data assumption that the source and target domains can have different output spaces, i.e., each domain can be associated with a different learning task (TL_survey_2010). In the context of transfer learning, it assumes that a generic nonlinear feature representation function is shared across all tasks. Then each task is associated with a shared representation learning function and a task-specific prediction function. In (task_diversity), task diversity is defined to characterize how the worst-case representation difference can be controlled when the task-averaged representation difference is small. In this case, the worst-case representation difference is the distance between two representation functions with the worst-case task-specific prediction function, while the task-averaged representation difference indicates the distance between two representation functions over all the training tasks.
Definition 3.3 (Task Diversity (task_diversity)).
@@
(14) $\underbrace{\sup_{h_{0}\in\mathcal{H}_{0}}\inf_{h^{\prime}\in\mathcal{H}}\left\{\mathcal{E}_{T}\left(h^{\prime}\circ g^{\prime}\right)-\mathcal{E}_{T}\left(h_{0}\circ g\right)\right\}}_{\text{Worst-case representation difference}}\leq\frac{1}{\nu}\cdot\underbrace{\left(\frac{1}{N}\sum_{i=1}^{N}\inf_{h^{\prime}\in\mathcal{H}}\left\{\mathcal{E}_{S_{i}}\left(h^{\prime}\circ g^{\prime}\right)-\mathcal{E}_{S_{i}}\left(h_{i}\circ g\right)\right\}\right)}_{\text{Task-averaged representation difference}}~~+~~~\epsilon$
@@
Based on the task diversity, (task_diversity) derive the excess risk bounds of transfer learning for the target task in terms of the complexity of the shared representation function class $\mathcal{G}$, the complexity of the prediction function class $\mathcal{H}$, the number of tasks $N$, and the number of training samples for each task (both source and target). Furthermore, (WatkinsUNA23) show that under the Lipschitz assumption for the loss function, the excess risk in the target task only achieves the standard rate of $\mathcal{O}(n_{T}^{-1/2})$, where $n_{T}$ is the number of training samples in the target task. By using the smoothness assumption for the loss function (srebro2010optimistic), they derive optimistic rates that interpolate between the standard rate of $\mathcal{O}(n_{T}^{-1/2})$ and the fast rate of $\mathcal{O}(n_{T}^{-1})$ for the excess risk in the target task.
@@
In addition, (du2021fewshot; tripuraneni2021provable) consider a simplified version of task diversity in the cases of linear prediction functions and quadratic loss. They also theoretically show the benefits (i.e., reduced sample complexity in the target task induced by all available source samples) of representation learning from source tasks. Based on the task diversity, (xu2021representation) analyze more realistic learning scenarios in which the source and target tasks use different prediction function spaces. However, all the aforementioned theoretical analyses assume uniform sampling from each source task, i.e., all source tasks are equally important for learning a representation function. Instead, (chen2022active; chen2023active; wang2023improved) study active transfer learning by quantifying the task relatedness and selecting the source tasks that are most relevant to the target task. Similarly, (xu2024towards) explore the selection of source tasks for multi-task fine-tuning of foundation models, e.g., fine-tuning the foundation model on auxiliary source tasks before adapting it to the target task with limited labeled samples. More recently, (zhao2023blessing) show that pre-training on a single source task with a high diversity of classes can provably improve the sample efficiency of the downstream tasks. In contrast, (cole2024provable) leverage task diversity to understand the in-context learning behavior of foundation models.
(a) Transferability estimation for selecting the source model
(b) Transferability estimation for selecting the target domain/task
Figure 3. Evaluation of transferability between the pre-trained source model and the target data: (a) Transferability scores select the best source model for the target data given a large pool of pre-trained source models. (b) Transferability scores identify the most suitable application domains/tasks for a source model.
3.1.3. Transferability Estimation
@@
In contrast to the data-centric transferability analyses in Subsection 3.1.1 and Subsection 3.1.2, this subsection explores the knowledge transferability of pre-trained source models. This is driven by the rapidly expanding open-source model repositories such as HuggingFace (Transformers) and PyTorch Hub (PyTorch). Fine-tuning a pre-trained source model on downstream target data sets with limited sample sizes improves model accuracy and robustness (HendrycksLM19). A natural question arises in this scenario: Given a large pool of pre-trained source models, how can we efficiently select the best one for a target data set? As shown in Fig. 3, another relevant question is how to identify the most suitable domains/tasks for a given pre-trained source model. One trivial solution is brute-force fine-tuning, where all source models are fine-tuned individually and then ranked based on their transfer accuracy. However, this method is highly time-consuming and computationally expensive. To solve this problem, transferability estimation has been studied to quantitatively measure how effectively the knowledge can be transferred from a pre-trained source model to a target domain (H-Score; NCE; LEEP; AgostinelliPUMF22; IbrahimPM22). Following (NCE), given a pre-trained source model $f_{S}(\cdot)$, the transferability from $f_{S}(\cdot)$ to a target domain associated with sampling distribution $\mathrm{P}_{T}$ can be defined below.
Definition 3.4 (Transferability Measure (NCE)).
@@
(15) $\mathrm{Trf}(S\to T)=\mathbb{E}_{(x,y)\sim\mathrm{P}_{T}}\left[\mathrm{acc}\left(x,y;f_{T}\right)\right]$
@@
The estimation of the transferability score is related to the architectures of pre-trained source models. In scenarios where the source models are trained on supervised classification tasks, every source model $f_{S}(\cdot)$ is associated with a feature extractor and a predictor. In this case, NCE (NCE) leverages conditional entropy to define the transferability score, assuming that source and target domains share the same input samples but different labels. It is motivated by the observation that the optimal average log-likelihood on target training samples is lower bounded by the negative conditional entropy. Similarly, LEEP (LEEP) estimates the average log-likelihood of target samples using the dummy label distributions generated from the pre-trained source model. Notably, it is shown that LEEP is an upper bound of the NCE (NCE) plus the average log-likelihood of the dummy labels. It is computationally efficient by using only a single forward pass of the source model through the target data. Nevertheless, LEEP can not handle unsupervised and self-supervised pre-trained models with only a feature extractor.
To address this problem, recent works (PARC; LogME; NguyenTHDTHN23) have utilized only the feature extractor $f_{S}(\cdot)$ to define transferability scores. They can be divided into two frameworks. One is to measure the class separability of the target samples in the feature space induced by $f_{S}(\cdot)$. For example, H-score (H-Score) is defined based on the inter-class variance and feature redundancy of target samples learned from the pre-trained source model. It is inspired by the connection between the optimal prediction error and the modal decomposition of the divergence transition matrix (huang2024universal). The follow-up work (IbrahimPM22) introduces a shrinkage-based H-score to improve the covariance estimation of H-score (H-Score) in high-dimensional feature spaces. $\mathcal{N}$LEEP (NLEEP) replaces the dummy label generation module of LEEP (LEEP) with a new Gaussian Mixture Model (GMM). Furthermore, GBC (GBC) maps each target class as a Gaussian distribution and estimates the pair-wise class separability (i.e., the amount of overlap between two class-wise Gaussian distributions) using the Bhattacharyya coefficient (bhattacharyya1946measure). The other one is to add a probabilistic linear transformation that maps the feature space of $f_{S}(\cdot)$ to the target output space in a Bayesian framework. For example, LogME (LogME) is defined over marginalized likelihood $p(y_{i}|x_{i};f_{S})=\int p(w)p(y_{i}|f_{S}(x_{i}),w)dw$, assuming that the prior distribution of the newly added linear transformation $w$ is an isotropic multivariate Gaussian $w\sim\mathcal{N}(0,\alpha^{-1}\mathbb{I})$. The follow-up work PACTran (PACTran) further defines a theoretically grounded family of transferability scores based on the optimal PAC-Bayesian error bound (GermainBLL16), taking into consideration various instantiations of the prior distribution for the linear transformation $w$, such as Dirichlet, Gamma, and Gaussian priors. Additionally, TransRate (TransRate) exploits the coding rate to estimate transferability scores of any intermediate layer within the pre-trained model. More recently, inspired by neural scaling laws for fine-tuned LLMs (tay2022scale), (lin2024selecting) investigatet the transferability estimation of large language models (LLMs) based on a rectified scaling law that characterizes the connection between the fine-tuned test loss and the number of target samples.
@@
The transferability metrics mentioned above enable the selection of the best source model from a large pool of open-sourced pre-trained models. Recent studies (SFDA; BBGNKB23) take one step further by studying source model ensemble selections, which transfer knowledge from multiple pre-trained source models to target training samples. This line of research is inspired by the success of ensemble machine learning models (Lakshminarayanan17) in improving model performance. Specifically, (AgostinelliUMF22) extend LEEP (LEEP) to select source model ensembles under the assumption that the source models operate independently. In contrast, OSBORN (BBGNKB23) relaxes this assumption and explores the inter-model cohesion among source models to estimate the transferability of an ensemble of models to a target domain.
3.2. Non-IID Transferability